OpenStudy (anonymous):

find the slope of the line by solving for y x-3y=-2 I got 2/3 but I don't know if it's right. Please check my answer

OpenStudy (anonymous):

-3y=-2-x y=(-2-x)/-3 y=(1/3)x+2/3 so the slope would be 1/3